### v2

# CarSpeedNet: A Deep Neural Network-based Car Speed Estimation from Smartphone Accelerometer

Abstract
We introduce the CarSpeedNet, a deep learning model designed to estimate car speed using three-axis accelerometer data from smartphones. Using 13 hours of data collected from a smartphone in cars across various roads, CarSpeedNet accurately models the relationship between smartphone acceleration and car speed. Ground truth speed data was collected at 1 [Hz] from GPS receivers. The model provides high-frequency speed estimation by incorporating historical data and achieves a precision of less than 0.72 [m/s] during extended driving tests, relying solely on smartphone accelerometer data without any connection to the car.
Index Terms:
Machine learning, supervised learning, deep neural network, long short-term memory, dead reckoning, car speed, inertial measurement unit, accelerometer.
I Introduction
In the domain of vehicular technology and traffic management, precise car speed estimation remains a fundamental challenge, integral to optimizing traffic systems and enhancing navigation reliability [1, 2, 3, 4, 5]. Generally, speed measurement has primarily relied on odometers and other embedded car sensors. However, these conventional methods are limited by their lack of real-time data-sharing capabilities, which is increasingly critical in modern traffic management systems [6]. The shift towards data-driven approaches is essential in an era where real-time information exchange is indispensable [7, 8, 9, 10, 11].
Smartphones, equipped with advanced sensors, including accelerometers, present a novel and accessible tool for speed estimation [12, 13]. The accelerometer’s ability to capture specific forces along three orthogonal axes at high frequencies, often surpassing 100 [Hz] and in some cases reaching up to 500 [Hz], positions it as a valuable resource for monitoring car dynamics [3]. This high-frequency measurement capability is critical for capturing the nuanced variations in acceleration necessary for accurate speed estimation across diverse driving conditions.
Using smartphones for speed estimation introduces significant challenges, primarily in converting raw accelerometer data into reliable speed estimates. This process necessitates sophisticated data processing techniques, an area where deep learning (DL) has demonstrated substantial potential [14]. DL, with its capacity to model complex data patterns, is particularly well-suited for interpreting the intricate relationships between accelerometer signals and car speed [15]. The intersection of DL and smartphone sensor technology represents a fertile area of research in modern transportation systems [16, 17, 18, 19, 20, 21, 22, 23].
This study presents a DL architecture, CarSpeedNet, specifically designed to estimate car speed using only smartphone accelerometer data. The primary innovation of this research lies in its exclusive focus on accelerometer data, foregoing the need for additional sensors such as gyroscopes, which are not universally available in all smartphones [1, 24, 25, 26]. By leveraging a widely available sensor, this approach ensures broad accessibility across diverse smartphone models.
This work emphasizes a DL-based approach, marking a significant departure from traditional speed estimation techniques [25]. The implementation of DL models allows for the capture of complex temporal and spatial patterns in acceleration data, enabling the model to achieve higher accuracy than conventional methods, all while relying on sensors readily available in consumer devices.
The methodology adopted in this work follows a two-phase process: the first phase involves the collection of a comprehensive dataset of acceleration signals from multiple smartphones; the second phase applies a supervised learning paradigm to train the DL model.
The key contributions of this work include the development of CarSpeedNet, a DL model for estimating car speed using smartphone accelerometer data in a real-time manner. The study demonstrates how DL models can accurately interpret complex accelerometer patterns for precise speed estimation, while also exploring the optimal data length needed for accurate predictions. Additionally, a robust data collection and training methodology was implemented, ensuring real-world applicability.
The paper is structured as follows: Section II covers the learning approach and introduces CarSpeedNet. Section III presents results, including error metrics and performance analysis. Section IV concludes the paper.
II Learning Approach
II-A Data Acquisition and Dataset Creation
A dataset was collected using a Samsung Galaxy Smartphone, with accelerometer data recorded at 500 Hz. The data, consisting of 13.2 hours of car operation across various terrains, was split into training (10.3 hours), validation (2.6 hours), and testing (0.5 hours) subsets.
The data aggregation was extensively conducted across various geographical terrains within the state of Israel, encompassing an array of road types, predominantly highways and urban streets. It is imperative to note that the data collection was strictly confined to on-road scenarios, deliberately excluding off-road data acquisition. Each driving session commenced only after attaining a stabilized GPS signal, ascertained through the Global Dilution of Precision (GDOP) metric, thereby minimizing locational inaccuracies [27, 4]. The raw data underwent meticulous processing, where the IMU data was subjected to low-pass filtering to mitigate noise interference and subsequently downsampled to a frequency of 20 [Hz]. Concurrently, the car speed data, ascertained from the GPS readings, was maintained at a consistent recording frequency of 1 [Hz], and served as the ground truth label.
II-B Loss Function
The following mean square error (MSE) loss function was minimized in the training stage
${\cal L}=\frac{1}{{2N}}\sum\limits_{j=1}^{N}{{{\left({{s}_{j}^{GT}-{{{\hat{s}}}_{j}}\left({{\bf{a}};{\bf{W}}}\right)}\right)}^{2}}},$ (1)
where $\frac{1}{{2N}}$ represents the normalization factor, where $N$ is the total number of examples in the batch. The ${s}_{j}^{GT}$ stands for the ground truth value for the $j^{th}$ example, and ${{{\hat{s}}}_{j}}\left({{\bf{a}};{\bf{W}}}\right)$ represents the predicted value by the network for the same example. The prediction is a function of $\bf a$, the input acceleration data, and $\bf W$, the learnable weights of the neural network.
II-C Models Exploration
We explored and optimized five DL models before finalizing CarSpeedNet.
  • DNN*: Adopted from previous work [1]. The model, originally designed for both accelerometer and gyroscope data, was retrained exclusively for accelerometer inputs, with 13,031 trainable parameters. The architecture comprises one long short-term memory (LSTM) layer [28], two bidirectional LSTM layers [29], and a single dense layer to output the estimated speed.
  • LSTM: A sequential model with LSTM and Batch Normalization layers, designed to capture temporal features, consisting of 17,181 trainable parameters.
  • WaveNet: Known for processing temporal data [30], this model uses convolutional layers with varying dilation rates and has 239,937 trainable parameter.
  • Bi-LSTM: Combines bidirectional and standard LSTM layers with dropout, containing 26,251 trainable parameters.
  • ResNet-inspired: Integrates ResNet blocks with Bi-LSTM layers [31], featuring 95,043 trainable parameters for capturing temporal dependencies.
We finalized the architecture, designated as CarSpeedNet, through iterative refinements based on insights gained from each model iteration.
II-D The CarSpeedNet
II-D1 Architecture
The CarSpeedNet architecture processes 80 sequential data points from 3-axis accelerometer inputs over 4 seconds (shown in Figure 1). It features bidirectional LSTM layers, including one with 100 units, capturing long-range dependencies in both directions. The following four LSTM layers further refine the temporal information. Batch Normalization is applied throughout to improve training stability, with reduced numbers of units (50, 20, 20, 20). After the LSTM layers, the model uses three Conv1D layers for feature extraction with 64, 64, and 32 filters each, a constant kernel size of 3, followed by a fully connected layer with 32 parameters, with the final layer outputting the speed estimation. A Rectified Linear Unit (ReLU) [32] activation function was applied between the layers.
The architectural framework of CarSpeedNet is distinguished by its bifurcated structure, systematically designed to address the dual aspects of feature extraction: temporal and spatial: initially, the architecture focuses on the extraction of temporal features. This phase is crucial for discerning the sequential dependencies and dynamics inherent in time-series data, such as the accelerometer readings. By effectively capturing these temporal relationships, the model lays the groundwork for a nuanced understanding of the car’s motion over time. Subsequently, the model shifts its focus to the extraction of spatial features from the temporally processed data. This stage is instrumental in identifying and isolating spatial patterns and correlations within the data, which are critical for comprehending the multi-dimensional nature of the accelerometer inputs.
Figure 1: The CarSpeedNet architecture.
II-D2 Training
The training of CarSpeedNet involved 178,169 trainable parameters, utilizing the Adam optimizer [33] with an initial learning rate of 0.001. An exponential decay schedule reduced the learning rate over 30,000 steps with a decay rate of 0.2 to enhance training efficiency. The maximum number of epochs was set as 200, and training was performed with a batch size of 32.
Figure 2: The suggested approach: smartphone accelerometer data is processed, including downsampling and batch aggregation. The prepared data is input into CarSpeedNet to estimate the car’s speed.
III Results and Discussion
This section evaluates the performance of CarSpeedNet using two error metrics: Root Mean Square Error (RMSE) and Mean Absolute Error (MAE). We compare CarSpeedNet with other models and explore how input size impacts performance, particularly for real-time applications.
III-A Error Metrics
We assess predictive accuracy using RMSE and MAE, both measured in m/s. RMSE captures large errors effectively, while MAE provides an average error magnitude. The RMSE and the MAE are computed as follows:
$RMSE=\sqrt{\frac{1}{N}\sum\limits_{j=1}^{N}{{{\left({s_{j}^{GT}-{\hat{s}_{j}}}\right)}^{2}}}},$ (2)
and
$MAE=\frac{1}{N}\sum\limits_{j=1}^{N}{\left|{s_{j}^{GT}-{\hat{s}_{j}}}\right|}.$ (3)
III-B Comparative Analysis of Model Performances
Figure 3 shows the RMSE and MAE for each model. All the results in this subsection are computed for 1[s] accelerometer input (20 measurements). CarSpeedNet delivers excellent accuracy with an RMSE of 2.9 [m/s] and an MAE of 1.3 [m/s], alongside a latency of 104.8 [ms]. The ResNet model follows closely with an RMSE of 2.9 [m/s] and an MAE of 1.4 [m/s]. Other models, like LSTM and WaveNet, show higher error rates and longer latencies, highlighting the superiority of CarSpeedNet in both accuracy and computational load.
Figure 3: Models’ performance evaluation using the RMSE and MAE metrics.
III-C Effect of input size on CarSpeedNet
We examined how varying input sizes (window length) impacted model accuracy and latency (Table I). Larger input window sizes (e.g., 4 [s]) improve accuracy (RMSE of 1.8 [m/s], MAE of 0.72 [m/s]) but introduce higher latency. CarSpeedNet’s ability to maintain high accuracy, especially at slower speeds, ensures reliable performance across diverse driving conditions, including detecting stationary states.
Additionally, the temporal performance of CarSpeedNet, encompassing the comparison between the estimated speeds by the model and the actual speeds (ground truth), was plotted over time for each input window size. These plots, presented in Figure 5, offer a visual representation of the model’s performance across different window sizes.
It is observed that CarSpeedNet demonstrates a high degree of accuracy in estimating speeds during slow-moving conditions. This precision at lower velocities is a testament to the model’s sensitivity and effectiveness in capturing subtle variations in the accelerometer data, which are prevalent at reduced speeds.
Significantly, the model’s proficiency is further highlighted in its capability to accurately detect stationary conditions. In instances where the car is at a complete halt, CarSpeedNet’s estimated speed closely aligns with zero. This ability to discern and accurately indicate a state of rest is a critical feature for any speed estimation model, as it ensures reliable performance across the entire spectrum of car motion, from high-speed travel to complete stops.
TABLE I: Input Size Effect
Window size RMSE [m/s] MAE [m/s] Latency [ms]
5 / 0.25[s] 4.9 2.5 96.5
10 / 0.5[s] 4.2 2.1 177
20 / 1[s] 2.9 1.3 100
40 / 2[s] 2.3 1 94
60 / 3[s] 2.2 0.88 102.4
80 / 4[s] 1.8 0.72 104.8
Figure 4: CarSpeedNet performance as a function of the input size.
Figure 5: CarSpeedNet vs. GT speed comparison, $w\in\{10,40,80\}$.
IV Conclusions
We introduced CarSpeedNet, a deep learning model designed for car speed estimation using smartphone accelerometer data only. The model achieves an error of less than 0.72 [m/s], outperforming traditional car-based systems and offering a universally accessible solution without requiring additional sensors. Larger input windows enhance accuracy, with the trade-off of increased latency. Also, the rigorous data collection and training methodology validate the model’s effectiveness in diverse driving conditions, including slow-speed and stationary states. CarSpeedNet provides a robust and scalable solution for intelligent transportation systems, opening opportunities for further applications in traffic management and safety, leveraging existing smartphone technology.