### v3

# CarSpeedNet: Learning-Based Speed Estimation from Accelerometer-Only Inertial Sensing

Abstract
Velocity estimation is a core component of state estimation and sensor fusion pipelines in mobile robotics and autonomous ground systems, directly affecting navigation accuracy, control stability, and operational safety. In conventional systems, velocity is obtained through wheel encoders, inertial navigation units, or tightly coupled multi-sensor fusion architectures. However, these sensing configurations are not always available or reliable, particularly in low-cost, redundancy-constrained, or degraded operational scenarios where sensors may fail, drift, or become temporarily unavailable.
This paper investigates the feasibility of estimating vehicle speed using only a single low-cost inertial sensor: a three-axis accelerometer embedded in a commodity smartphone. We present CarSpeedNet, a learning-based inertial estimation framework designed to infer speed directly from raw accelerometer measurements, without access to gyroscopes, wheel odometry, vehicle bus data, or external positioning during inference. From a sensor fusion perspective, this setting represents an extreme case of sensing sparsity, in which classical integration-based or filter-based approaches become unstable due to bias accumulation and partial observability.
Rather than explicitly estimating physical states such as orientation or sensor bias, the proposed approach performs implicit latent-state approximation from temporal accelerometer data. By exploiting temporal context, CarSpeedNet captures motion dynamics that are not observable instantaneously, enabling velocity inference under severe sensing constraints. Using real-world driving data collected across diverse environments, we demonstrate that the proposed method achieves accurate and robust speed estimation despite the absence of complementary sensors.
The results are relevant to mobile robots, autonomous ground vehicles, and low-cost navigation systems, where robustness to sensor degradation and minimal sensing requirements are critical design considerations.
I Introduction
Accurate velocity estimation is a core requirement in mobile robotics, autonomous ground vehicles, and intelligent sensing platforms. Speed information directly influences dead reckoning, motion control, localization, and higher-level decision-making. In conventional robotic and vehicular systems, velocity is typically obtained through wheel encoders, inertial navigation systems (INS), or tightly coupled sensor fusion frameworks that integrate inertial and exteroceptive measurements [farrell2008aided, or2023learning]. However, such sensing configurations are not universally available and may suffer from degradation due to wheel slippage, sensor bias, calibration drift, or hardware cost constraints [kerber2018data].
In parallel, the proliferation of smartphones has introduced a class of low-cost inertial sensing platforms. However, these devices present unique challenges: unknown mounting orientations, irregular sampling jitters, and uncalibrated sensor biases[cite: 18]. From a robotics perspective, treating a smartphone as a proxy Inertial Measurement Unit (IMU) requires models that are robust to these stochastic perturbations [mantouka2021smartphone]. The accelerometer’s ability to capture high-frequency specific forces makes it a viable candidate for motion inference if the gravity-acceleration entanglement can be resolved through learned latent representations [or2023learning].
Estimating velocity using only accelerometer data is inherently challenging, as classical inertial navigation approaches based on direct integration are highly sensitive to sensor bias, noise accumulation, and step-size selection, leading to rapid drift and unreliable estimates in practical deployments [farrell2008aided, or2022adaptive, sharp2009gdop]. The difficulty of this problem is fundamentally tied to observability. Linear acceleration alone does not uniquely determine velocity without complementary measurements or prior assumptions. Nevertheless, recent advances in deep learning (DL) and adaptive learning mechanisms have demonstrated the potential of data-driven models to capture implicit temporal and dynamical patterns that are difficult to model analytically, and to mitigate some of the limitations inherent to classical inertial estimation pipelines [lecun2015deep, li2021inertial, or2022adaptive]. In particular, learning-based approaches can exploit correlations across time, motion regimes, and repeated driving or motion patterns, enabling partial compensation for missing sensors in constrained sensing setups [freydin2022learning, wang2017deepspeedometer, abdelgawad2019estimating].
Within this context, this paper investigates learning-based velocity estimation under strict sensor minimalism. We introduce CarSpeedNet, a deep neural network architecture designed to estimate speed using only three-axis accelerometer measurements acquired from a smartphone. The proposed approach operates without gyroscope data, wheel odometry, vehicle communication interfaces, or external positioning systems during inference. This setting reflects a highly constrained yet practically relevant scenario, analogous to low-cost mobile robots, redundant sensing layers in autonomous systems, or temporary navigation modules deployed in environments with limited sensing infrastructure [campolo2012smartcar, astarita2018mobile, ali2021traffic].
Rather than focusing solely on architectural novelty, this work emphasizes system-level insight. In particular, we examine how temporal window length affects estimation accuracy and latency, illustrating the trade-off between information accumulation and real-time responsiveness. Such trade-offs are central to robotic sensing pipelines, where estimation performance must be balanced against computational load and control-loop requirements [tong2022vehicle]. The results demonstrate that longer temporal contexts improve robustness and accuracy, while shorter windows favor low-latency applications, offering practical guidance for deployment under different operational constraints. The contributions of this work are threefold. First, we demonstrate the feasibility of accelerometer-only velocity estimation using a learning-based approach under real-world conditions. Second, we provide empirical and conceptual insight into the role of temporal context in learning-based inertial speed estimation, interpreting the temporal window as an information horizon under partial observability. Third, we frame accelerometer-only velocity estimation as a latent-state approximation problem under sensor minimalism, offering practical guidance for robotic and autonomous systems operating under limited sensing capabilities.
The remainder of the paper is organized as follows. Section II presents the learning approach and the CarSpeedNet architecture. Section III provides a conceptual discussion of learning-based velocity estimation under sensor minimalism, including partial observability and latent-state approximation. Section IV reports experimental results and discusses system-level trade-offs. Section V concludes the paper and outlines directions for future research.
II Learning Approach
II-A Data Acquisition and Dataset Creation
A dataset was collected using a Samsung Galaxy Smartphone, with accelerometer data recorded at 500 Hz. The data, consisting of 13.2 hours of car operation across various terrains, was split into training (10.3 hours), validation (2.6 hours), and testing (0.5 hours) subsets.
The data aggregation was extensively conducted across various geographical terrains within the state of Israel, encompassing an array of road types, predominantly highways and urban streets. It is imperative to note that the data collection was strictly confined to on-road scenarios, deliberately excluding off-road data acquisition. Each driving session commenced only after attaining a stabilized GPS signal, ascertained through the Global Dilution of Precision (GDOP) metric, thereby minimizing locational inaccuracies [sharp2009gdop, farrell2008aided]. The raw data underwent meticulous processing, where the IMU data was subjected to low-pass filtering to mitigate noise interference and subsequently downsampled to a frequency of 20 [Hz]. Concurrently, the car speed data, ascertained from the GPS readings, was maintained at a consistent recording frequency of 1 [Hz], and served as the ground truth label.
II-B Loss Function
The following mean square error (MSE) loss function was minimized in the training stage
${\cal L}=\frac{1}{{2N}}\sum\limits_{j=1}^{N}{{{\left({{s}_{j}^{GT}-{{{\hat{s}}}_{j}}\left({{\bf{a}};{\bf{W}}}\right)}\right)}^{2}}},$ (1)
where $\frac{1}{{2N}}$ represents the normalization factor, where $N$ is the total number of examples in the batch. The ${s}_{j}^{GT}$ stands for the ground truth value for the $j^{th}$ example, and ${{{\hat{s}}}_{j}}\left({{\bf{a}};{\bf{W}}}\right)$ represents the predicted value by the network for the same example. The prediction is a function of $\bf a$, the input acceleration data, and $\bf W$, the learnable weights of the neural network.
II-C Models Exploration
We explored and optimized five DL models before finalizing CarSpeedNet.
  • DNN*: Adopted from previous work [freydin2022learning]. The model, originally designed for both accelerometer and gyroscope data, was retrained exclusively for accelerometer inputs, with 13,031 trainable parameters. The architecture comprises one long short-term memory (LSTM) layer [hochreiter1997long], two bidirectional LSTM layers [zhou2016attention], and a single dense layer to output the estimated speed.
  • LSTM: A sequential model with LSTM and Batch Normalization layers, designed to capture temporal features, consisting of 17,181 trainable parameters.
  • WaveNet: Known for processing temporal data [oord2016wavenet], this model uses convolutional layers with varying dilation rates and has 239,937 trainable parameter.
  • Bi-LSTM: Combines bidirectional and standard LSTM layers with dropout, containing 26,251 trainable parameters.
  • ResNet-inspired: Integrates ResNet blocks with Bi-LSTM layers [he2016deep], featuring 95,043 trainable parameters for capturing temporal dependencies.
These architectures were selected to represent a range of temporal modeling strategies commonly used in inertial and time-series learning. We finalized the architecture, designated as CarSpeedNet, through iterative refinements based on insights gained from each model iteration.
II-D The CarSpeedNet
II-D1 Architecture
The CarSpeedNet architecture processes 80 sequential data points from 3-axis accelerometer inputs over 4 seconds (shown in Figure 1). It features bidirectional LSTM layers, including one with 100 units, capturing long-range dependencies in both directions. The following four LSTM layers further refine the temporal information. Batch Normalization is applied throughout to improve training stability, with reduced numbers of units (50, 20, 20, 20). After the LSTM layers, the model uses three Conv1D layers for feature extraction with 64, 64, and 32 filters each, a constant kernel size of 3, followed by a fully connected layer with 32 parameters, with the final layer outputting the speed estimation. A Rectified Linear Unit (ReLU) [nair2010rectified] activation function was applied between the layers.
The architectural framework of CarSpeedNet is distinguished by its bifurcated structure, systematically designed to address the dual aspects of feature extraction: temporal and spatial: initially, the architecture focuses on the extraction of temporal features. This phase is crucial for discerning the sequential dependencies and dynamics inherent in time-series data, such as the accelerometer readings. By effectively capturing these temporal relationships, the model lays the groundwork for a nuanced understanding of the car’s motion over time. Subsequently, the model shifts its focus to the extraction of spatial features from the temporally processed data. This stage is instrumental in identifying and isolating spatial patterns and correlations within the data, which are critical for comprehending the multi-dimensional nature of the accelerometer inputs.
Figure 1: The CarSpeedNet architecture.
II-D2 Training
The training of CarSpeedNet involved 178,169 trainable parameters, utilizing the Adam optimizer [kingma2014adam] with an initial learning rate of 0.001. An exponential decay schedule reduced the learning rate over 30,000 steps with a decay rate of 0.2 to enhance training efficiency. The maximum number of epochs was set as 200, and training was performed with a batch size of 32.
Figure 2: The suggested approach: smartphone accelerometer data is processed, including downsampling and batch aggregation. The prepared data is input into CarSpeedNet to estimate the car’s speed.
III Conceptual Discussion: Learning Under Sensor Minimalism
This section provides a system-level interpretation of accelerometer-only velocity estimation and reframes the proposed approach as a latent-state approximation problem under partial observability. The discussion is intended to complement the empirical results by clarifying why learning-based models can succeed where classical estimation pipelines fail, and how temporal window length governs the available information for inference.
III-A Partial Observability in Accelerometer-Only Velocity Estimation
Under accelerometer-only sensing, the velocity state is not directly observable. Linear acceleration measurements alone do not uniquely determine velocity in the presence of unknown orientation, sensor bias, and measurement noise. Multiple velocity trajectories may generate indistinguishable accelerometer signals over short time intervals, particularly during constant-velocity or low-dynamic motion regimes. As a result, the instantaneous mapping from acceleration to velocity is inherently non-injective, rendering the estimation problem partially observable. Figure 3 illustrates this ambiguity: distinct latent velocity trajectories can produce locally indistinguishable accelerometer measurements within a finite temporal window, making velocity inference ill-posed without additional context.
III-B Limitations of Classical Estimation Approaches
Classical inertial navigation methods rely on explicit integration of acceleration measurements, often augmented with bias estimation and orientation tracking. In accelerometer-only settings, these assumptions break down: small biases accumulate rapidly, orientation uncertainty contaminates projected acceleration, and integration amplifies noise over time. Without external corrections or adaptive mechanisms, classical estimators diverge quickly, making them unsuitable for sustained velocity estimation under strict sensor minimalism.
III-C Learning as Latent-State Approximation
In contrast to classical estimators, learning-based models do not explicitly estimate physical states such as sensor bias or orientation. Instead, they operate as latent-state approximators. By observing sequences of accelerometer measurements, a neural network can implicitly encode latent information about motion dynamics, bias patterns, and regime transitions that are not directly observable at any single time instant. This enables approximate velocity inference even when the underlying physical system is only partially observable.
Formally, the proposed approach estimates velocity using a learned mapping
$\hat{v}_{t}=f_{\theta}\!\left(a_{t-W+1:t}\right),$ (2)
where $a_{t-W+1:t}$ denotes a temporal window of accelerometer measurements of length $W$, and $f_{\theta}(\cdot)$ represents a neural network parameterized by $\theta$. While instantaneous acceleration does not uniquely define velocity, the extended temporal context allows the model to disambiguate motion regimes and approximate latent state information.
III-D Temporal Window Length as an Information Horizon
From this perspective, the temporal window length $W$ acts as an information horizon. Short windows provide limited contextual information and are insufficient to resolve ambiguities inherent to partial observability. Increasing $W$ expands the available motion history, enabling the model to accumulate evidence across stationary, constant-velocity, and dynamic regimes. This reduces uncertainty in velocity inference at the cost of increased latency. The empirical results in Section III can therefore be interpreted as a study of the trade-off between information accumulation and real-time responsiveness under sensor-minimal conditions.
Latent velocityMeasured accel.trajectory 1trajectory 2observablelong window
Figure 3: Latent-state ambiguity under accelerometer-only sensing. Distinct latent velocity trajectories can produce locally indistinguishable accelerometer measurements. The shaded region denotes a temporal window; longer windows provide sufficient context for learning-based latent-state approximation under partial observability.
IV Results and Discussion
This section evaluates the performance of CarSpeedNet using two error metrics: Root Mean Square Error (RMSE) and Mean Absolute Error (MAE). We compare CarSpeedNet with other models and explore how input size impacts performance, particularly for real-time applications.
IV-A Error Metrics
We assess predictive accuracy using RMSE and MAE, both measured in m/s. RMSE captures large errors effectively, while MAE provides an average error magnitude. The RMSE and the MAE are computed as follows:
$RMSE=\sqrt{\frac{1}{N}\sum\limits_{j=1}^{N}{{{\left({s_{j}^{GT}-{\hat{s}_{j}}}\right)}^{2}}}},$ (3)
and
$MAE=\frac{1}{N}\sum\limits_{j=1}^{N}{\left|{s_{j}^{GT}-{\hat{s}_{j}}}\right|}.$ (4)
IV-B Comparative Analysis of Model Performances
Figure 4 shows the RMSE and MAE for each model. All the results in this subsection are computed for 1[s] accelerometer input (20 measurements). CarSpeedNet achieves an RMSE of 2.9[m/s][m/s] and an MAE of 1.3[m/s][m/s] for a 1-second window, outperforming the WaveNet and Bi-LSTM baselines by over 40% in error reduction, alongside a latency of 104.8 [ms]. The ResNet model follows closely with an RMSE of 2.9 [m/s] and an MAE of 1.4 [m/s]. Other models, like LSTM and WaveNet, show higher error rates and longer latencies, highlighting the superiority of CarSpeedNet in both accuracy and computational load.
Figure 4: Models’ performance evaluation using the RMSE and MAE metrics.
IV-C Effect of input size on CarSpeedNet
We examined how varying input sizes (window length) impacted model accuracy and latency (Table I). Larger input window sizes (e.g., 4 [s]) improve accuracy (RMSE of 1.8 [m/s], MAE of 0.72 [m/s]) but introduce higher latency. CarSpeedNet’s ability to maintain high accuracy, especially at slower speeds, ensures reliable performance across diverse driving conditions, including detecting stationary states. These results support the interpretation of the temporal window as an information horizon, where increased window length improves latent-state approximation under partial observability at the cost of increased inference latency, as conceptually illustrated in Fig. 3.
Additionally, the temporal performance of CarSpeedNet, encompassing the comparison between the estimated speeds by the model and the actual speeds (ground truth), was plotted over time for each input window size. These plots, presented in Figure 5, offer a visual representation of the model’s performance across different window sizes.
It is observed that CarSpeedNet demonstrates a high degree of accuracy in estimating speeds during slow-moving conditions. This precision at lower velocities is a testament to the model’s sensitivity and effectiveness in capturing subtle variations in the accelerometer data, which are prevalent at reduced speeds.
Significantly, the model’s proficiency is further highlighted in its capability to accurately detect stationary conditions. In instances where the car is at a complete halt, CarSpeedNet’s estimated speed closely aligns with zero. This ability to discern and accurately indicate a state of rest is a critical feature for any speed estimation model, as it ensures reliable performance across the entire spectrum of car motion, from high-speed travel to complete stops.
TABLE I: Input Size Effect
Window size RMSE [m/s] MAE [m/s] Latency [ms]
5 / 0.25[s] 4.9 2.5 96.5
10 / 0.5[s] 4.2 2.1 177
20 / 1[s] 2.9 1.3 100
40 / 2[s] 2.3 1 94
60 / 3[s] 2.2 0.88 102.4
80 / 4[s] 1.8 0.72 104.8
Figure 5: CarSpeedNet performance as a function of the input size.
Figure 6: CarSpeedNet vs. GT speed comparison, $w\in\{10,40,80\}$.
V Conclusions
We introduced CarSpeedNet, a deep learning model designed for car speed estimation using smartphone accelerometer data only. The model achieves an error of less than 0.72 [m/s], achieving competitive accuracy relative to traditional car-based sensing approaches. Larger input windows enhance accuracy, with the trade-off of increased latency. Also, the rigorous data collection and training methodology validate the model’s effectiveness in diverse driving conditions, including slow-speed and stationary states. CarSpeedNet provides a learning-based latent-state approximation mechanism for velocity estimation under sensor-minimal conditions, opening opportunities for further applications in traffic management and safety, leveraging existing smartphone technology. From a robotics perspective, the proposed approach highlights the potential of learning-based velocity estimation under severe sensing constraints. Accelerometer-only speed inference is relevant to low-cost mobile robots, auxiliary sensing modules, and redundant navigation layers in autonomous systems where wheel odometry, gyroscopes, or external positioning may be unavailable or unreliable. While not intended to replace full sensor fusion pipelines, CarSpeedNet demonstrates how data-driven models can complement classical estimation methods in constrained operational settings. Future work will investigate the integration of CarSpeedNet as a redundant safety layer in industrial AGVs to detect wheel-slip conditions where classical encoders fail.
[Uncaptioned image] Barak Or (Member, IEEE) received a B.Sc. degree in aerospace engineering (2016), a B.A. degree (cum laude) in economics and management (2016), and an M.Sc. degree in aerospace engineering (2018) from the Technion–Israel Institute of Technology. He graduated with a Ph.D. degree from the University of Haifa, Haifa (2022). His research interests include navigation, deep learning, sensor fusion, and estimation theory.